# algorithmic

### v1

@@ Abstract
Choosing a meaningful subset of features from high-dimensional observations in unsupervised settings can greatly enhance the accuracy of downstream analysis, such as clustering or dimensionality reduction, and provide valuable insights into the sources of heterogeneity in a given dataset. In this paper, we propose a self-supervised graph-based approach for unsupervised feature selection. Our method’s core involves computing robust pseudo-labels by applying simple processing steps to the graph Laplacian’s eigenvectors. The subset of eigenvectors used for computing pseudo-labels is chosen based on a model stability criterion. We then measure the importance of each feature by training a surrogate model to predict the pseudo-labels from the observations. Our approach is shown to be robust to challenging scenarios, such as the presence of outliers and complex substructures. We demonstrate the effectiveness of our method through experiments on real-world datasets, showing its robustness across multiple domains, particularly its effectiveness on biological datasets.
@@ 1 Introduction
Improvements in sampling technology enable scientists across many disciplines to acquire numerous variables from biological or physical systems. One of the critical challenges in real-world scientific data is the presence of noisy, information-poor, or nuisance features. While such features could be mildly harmful to supervised learning, they could dramatically affect the outcome of downstream analysis tasks (e.g., clustering or manifold learning) in the unsupervised setting [Mahdavi et al., 2019]. There is thus a growing need for unsupervised feature selection schemes that enhance latent signals of interest by removing nuisance variables and thus advance reliable data-driven scientific discovery.
Unsupervised Feature Selection (UFS) methods are designed to identify a set of informative features that can improve the outcome of downstream analysis tasks such as clustering and manifold learning. With the lack of labels, however, selecting features becomes a challenge since the downstream task cannot be used to drive the selection of features. As an alternative, most UFS methods use a label-free criterion that correlates with the downstream task. For instance, many UFS schemes rely on a reconstruction prior [Li et al., 2017] and seek a subset of features that can be used to reconstruct the entire set of features as accurately as possible. Several works use Autoencoders (AE) to learn a reduced representation of the data while applying a sparsification penalty to force the AE to remove redundant features. This idea was implemented with several types of sparsity-inducing regularizers, including $\ell_{2,1}$ based [Chandra and Sharma, 2015, Han et al., 2018], relaxed $\ell_{0}$ [Balın et al., 2019, Shaham et al., 2022, Svirsky and Lindenbaum, ] and more.
One of the most commonly used criteria for UFS is feature smoothness. According to this hypothesis, the structure of interest, such as clusters or a manifold, can be captured using the graph Laplacian matrix [Ng et al., 2001]. The smoothness of features is measured using the Laplacian Score (LS) [He et al., 2005], which is based on the Rayleigh quotient of the Laplacian. A feature that is smooth with respect to the graph is considered to be associated with the primary underlying data structures. There are many other UFS methods that use a graph to select informative features Li et al. [2018], Roffo et al. [2017], Zhu et al. [2017, 2020], Xie et al. [2023]. [Li et al., 2012] derived Nonnegative Discriminative Feature Selection (NDFS) , which performs feature selection and spectral clustering simultaneously. Its extension Li and Tang [2015] adds a loss term to prevent the joint selection of correlated features.
@@
In this work, we present a spectral self-supervised scheme for feature selection. The key idea is to selectively and discriminatively use the eigenvectors of the graph Laplacian. We implement this process through a multi-stage approach. Firstly, we generate robust discrete pseudo-labels from the eigenvectors and filter them based on a stability measure. Next, we fit flexible surrogate classification models on the selected eigenvectors and query the models for feature scores. Using these components, we can identify informative features that are effective for clustering on real-world datasets.
Figure 1: Illustration of SSFS. In (a) we show a tSNE scatter plot of noisy MNIST digits (3, 6, 8). (b) Presents the six leading eigenvectors computed based on the graph Laplacian of the data. Samples are ordered according to the identity of the digit. (c) We then use the $k$-medoids algorithm to define pseudo-labels ${\bm{y}}_{i}^{*}$. These are presented as colors overlayed on the eigenvectors. (d) We select the three eigenvectors whose pseud-labels are the most “stable” with respect to several prediction models (see Section 3.2). (e) For each data feature we estimate its importance score for each of the selected eigenvectors (see Section 3.3). (f) We aggregate the feature scores across eigenvectors.
@@ 2.1 Laplacian score and representation-based feature selection
A potential drawback of the Laplacian score is its dependence on many eigenvectors. This may reduce its stability in measuring a feature’s importance to the data’s main structures. To overcome this limitation, Zhao and Liu [2007] derived an alternative score based only on a feature’s correlation to the leading Laplacian eigenvectors. A related, more sophisticated approach is Multi-Cluster Feature Selection (MCFS) [Cai et al., 2010], which computes the solutions to the generalized eigenvector problem ${\bm{L}}{\bm{v}}=\lambda{\bm{D}}{\bm{v}}$. The leading eigenvectors are then used as pseudo-labels for a regression task with $l_{1}$ regularization. Specifically, MCFS applies Least Angle Regression (LARS) [Efron et al., 2004] to obtain, for each leading eigenvector ${\bm{v}}_{i}$, a sparse vector of coefficients $\bm{\beta}^{i}\in\mathbb{R}^{p}$. A feature score is computed by maximizing the absolute values of its corresponding coefficient, $s_{j}=\max_{i}|\beta^{i}_{j}|.$ The output of MCFS is the set of features with the highest score. In the next section, we derive Spectral Self-supervised Feature Selection (SSFS), which improves upon the MCFS algorithm in several critical aspects.
@@ 3.1 Rationale
where $|A_{i}|$ denotes the size of cluster $A_{i}$. In this scenario, the zero eigenvalue of the graph Laplacian has multiplicity $k$, and the corresponding eigenvectors are equal, up to a rotation matrix, to a matrix $E\in\mathbb{R}^{n\times d}$ whose columns are equal to ${\bm{e}}^{1},\ldots,{\bm{e}}^{k}$. In such a case, the $k$ leading eigenvectors are indeed suitable for use as pseudo-labels for the feature selection task. Assuming that the clusters are amenable to a linear separation, the MCFS algorithm should provide highly informative features in terms of cluster separation.
However, cluster separation is often imperfect in many applications, which can make using leading eigenvectors for regression suboptimal. Here are some common scenarios: 1) High-dimensional datasets may contain substructures in top eigenvectors, while the main structure of interest will appear later in spectrum. For illustration, consider the MNIST dataset visualized via t-SNE in Figure 1(a). The data contains images of $3,6$ and $8$. Panel (b) shows the elements of the six leading eigenvectors of the graph Laplacian matrix, sorted by their corresponding digits. The leading eigenvector shows a clear gap between images of digit $6$ and the rest of the data. However, there is no clear separation between digits $3$ and $8$. Indeed, the next eigenvector is not associated with such a separation. Applying feature selection with this eigenvector may produce spurious features irrelevant to separating the two digits. This scenario is prevalent in the real datasets used in the experimental section. For example, Figure 2a shows four eigenvectors of a graph computed from observations containing the genetic expression data from prostate cancer patients and controls [Singh et al., 2002]. The leading two eigenvectors, however, are not associated with the patient-control separation.
@@
Figure 2: The first four Laplacian eigenvectors of two real datasets. Samples are sorted according to the real class label and colored by the outcome of a one-dimensional $k$-medoids per eigenvector. The vertical bar indicates the separation between the classes. In Prostate-GE, ${\bm{v}}_{4}$ is the most informative to the class labels, and an outlier can be seen on the upper left in the third and fourth eigenvectors. In TOX-171, ${\bm{v}}_{3}$ is more informative to the class labels than ${\bm{v}}_{2}$.
@@ 3.4 Choice of Surrogate Models
8:     end for
9:     for $m=1$ to $p$ do
10:        Estimate the variance of the $m$-th feature score:
@@
11:     end for
12:     $\mathcal{\hat{S}}_{i}=\sum_{m=1}^{p}\widehat{\mathrm{Var}}({s}_{m}(h_{i}))$
13:     $\mathcal{\hat{S}}\leftarrow\mathcal{\hat{S}}\cup\{\mathcal{\hat{S}}_{i}$
14:  end for
15:  Select the indices of the $k$ smallest elements in $\mathcal{\hat{S}}$ and store in $I$
16:  return  $\mathcal{Y}^{*}$, $I$
@@ 4 The importance of a proper selection of eigenvectors: analysis of the product manifold model
As described in Section 3.1, the principle of selecting the leading Laplacian eigenvectors as pseudo-labels is inspired by the case of highly separable clusters, where observations in different clusters have very low connectivity between them. In many cases, the separation between meaningful states (i.e., biological or medical conditions) may not be that clear. To illustrate this point, consider the MNIST example in Figure 1. The separation between digit $6$ and the rest of the data is clear and appears in the leading Laplacian eigenvector. In contrast, digits $8$ and $3$ are not clearly separated. Figure 3a shows a scatter plot of these digits, where each image is located according to its coordinates in the third and fourth eigenvectors. Even when considering the most relevant eigenvectors, there is no clear separation between the digits. Instead, the transition between $3$ and $8$ is smooth and depends on the properties of the digits.
To provide some insight into the importance of eigenvector selection, we analyze a product of manifold model. Our analysis is based on results from two research topics: (i) the convergence, under the manifold assumption, of the Laplacian eigenvectors to the eigenfunctions of the Laplace Beltrami operator associated with the manifold, and (ii) the properties of manifold products. We next provide a brief background on these two topics.
@@ 4.1 Convergence of the Laplacian eigenvectors
In many applications, the high dimensional observations are assumed to reside close to some manifold $\mathcal{M}$ with low intrinsic dimensionality, which we denote by $d$. Many papers in recent decades have analyzed the relation between the Laplacian eigenvectors and the manifold structure Von Luxburg et al. [2008], Singer and Wu [2017], García Trillos et al. [2020], Wormell and Reich [2021], Dunson et al. [2021], Calder and Trillos [2022]. More formally, let ${\bm{v}}_{k}$ denote the $k$-th eigenvector of the graph Laplacian, and let $g_{k}$ denote the $k$-th eigenfunction of the Laplace-Beltrami (LB) operator. We usually assume that $g_{k}$ is normalized such that
@@
${\bm{v}}_{k}\xrightarrow[n\to\infty]{}g_{k}(X).$
where $g_{k}(X)$ is a vector of size $n$ containing samples of the function $g_{k}$ at the $n$ rows of the data matrix $X$.
Let us provide a simple example. Consider $n$ points sampled uniformly at random over an interval $[0,1]$. The LB operator over an interval is the second derivative whose eigenfunctions are the harmonic functions $g_{k}(x)=\cos(k\pi x)$. Figure 3 shows the three Laplacian eigenvectors computed with $n=10^{2},10^{3}$ and $3\cdot 10^{3}$ points. As $n\to\infty$, the difference between $g_{k}(X)$ and $v_{k}$ is decreases to 00.
Here, we use a convergence result from Cheng and Wu [2022], derived under the following assumptions: (i) The $n$ observations were generated according to a uniform distribution over the manifold, such that $\mu({\bm{x}})$ equals to a constant $\mu$. (ii) Let $\lambda_{k}$ denote the eigenvalue associated with the eigenfunction $g_{k}$. To ensure the stability of the eigenvectors, we assume a spectral gap between the smallest $K$ eigenvalues bounded away from 00 such that,
@@
(iii) The graph weights are computed by a Gaussian kernel $\exp(-\|x_{i}-x_{j}\|^{2}/\epsilon_{n})$, with a bandwidth $\epsilon_{n}\xrightarrow[n\to\infty]{}0^{+}$ that satisfies $\epsilon_{n}^{d/2+2}>C_{k}\frac{\log n}{n}$ for a constant $C_{K}$.
@@ 4.2 The product of manifold model
For simplicity, we denote by ${\bm{v}}_{l,k}$ the $(l,k)$-th eigenvector of the Laplacian matrix, as ordered by $\lambda_{l,k}$. An example of a product of $2$ manifolds is illustrated in Figure 4b. The figure shows the leading eight eigenvectors of the graph Laplacian. The eigenvectors are indexed by the vector ${\bm{b}}=[l,k]$. The full details of this example will be provided in the next section.
@@
Figure 3: Panel (a) shows a scatter plot of the noisy MNIST dataset, containing digits 3 and 8, where each image is located according to its coordinates in the third and fourth eigenvectors. Panel (b) shows the leading eigenvector of a graph computed over $n$ points on a $1D$ interval and the leading eigenfunction $\cos(\pi x)$.
4.3 Considerations for eigenvector selection in a product-of-manifold model

### v2

@@ Abstract
Choosing a meaningful subset of features from high-dimensional observations in unsupervised settings can greatly enhance the accuracy of downstream analysis tasks, such as clustering or dimensionality reduction, and provide valuable insights into the sources of heterogeneity in a given dataset. In this paper, we propose a self-supervised graph-based approach for unsupervised feature selection. Our method’s core involves computing robust pseudo-labels by applying simple processing steps to the graph Laplacian’s eigenvectors. The subset of eigenvectors used for computing pseudo-labels is chosen based on a model stability criterion. We then measure the importance of each feature by training a surrogate model to predict the pseudo-labels from the observations. Our approach is shown to be robust to challenging scenarios, such as the presence of outliers and complex substructures. We demonstrate the effectiveness of our method through experiments on real-world datasets from multiple domains, with a particular emphasis on biological datasets.
@@ 1 Introduction
Improvements in sampling technology enable scientists across many disciplines to acquire numerous variables from biological or physical systems. One of the critical challenges in real-world scientific data is the presence of noisy, information-poor, or nuisance features. While such features could be mildly harmful to supervised learning, they could dramatically affect the outcome of downstream analysis tasks (e.g., clustering or manifold learning) in unsupervised settings [Mahdavi et al., 2019]. There is thus a growing need for unsupervised feature selection schemes that enhance latent signals of interest by removing nuisance variables and thus advance reliable data-driven scientific discovery.
Unsupervised Feature Selection (UFS) methods are designed to identify a set of informative features that can improve the outcome of downstream analysis tasks such as clustering and manifold learning. With the lack of labels, however, selecting features becomes highly challenging since the downstream task cannot be used to drive the selection of features. As an alternative, most UFS methods use a label-free criterion that correlates with the downstream task. For instance, many UFS schemes rely on a reconstruction prior [Li et al., 2017] and seek a subset of features that can be used to reconstruct the entire dataset as accurately as possible. Several works use Autoencoders (AE) to learn a reduced representation of the data while applying a sparsification penalty to force the AE to remove redundant features. This idea was implemented with several types of sparsity-inducing regularizers, including $\ell_{2,1}$ based [Chandra and Sharma, 2015, Han et al., 2018], relaxed $\ell_{0}$ [Balın et al., 2019, Shaham et al., 2022, Svirsky and Lindenbaum, ] and more.
One of the most commonly used criteria for UFS is feature smoothness. A common hypothesis is that the structure of interest, such as clusters or a manifold, can be captured using a graph representation[Ng et al., 2001]. The smoothness of features is measured using the Laplacian Score (LS) [He et al., 2005], which is equal to the Rayleigh quotient of a normalized feature vector and the graph Laplacian. A feature that is smooth with respect to the graph is considered to be associated with the primary underlying data structures. There are many other UFS methods that use a graph to select informative features Li et al. [2018], Roffo et al. [2017], Zhu et al. [2017, 2020], Xie et al. [2023]. [Li et al., 2012] derived Nonnegative Discriminative Feature Selection (NDFS), which performs feature selection and spectral clustering simultaneously. Its extension Li and Tang [2015] adds a loss term to prevent the joint selection of correlated features.
@@
In this work, we present a spectral self-supervised scheme for feature selection. Our main innovation is to properly select a subset of Laplacian eigenvectors by a multi-stage approach. Firstly, we generate robust discrete pseudo-labels from the eigenvectors and filter them based on a stability measure. Next, we fit flexible surrogate classification models on the selected eigenvectors and query the models for feature scores. Using these components, we can identify informative features that are effective for clustering on real-world datasets.
Figure 1: Illustration of SSFS. Panel (a) shows a tSNE scatter plot of noisy MNIST digits (3, 6, 8). Panel (b) presents the six leading eigenvectors computed based on the graph Laplacian of the data. Samples are ordered according to the identity of the digit. (c) We apply the $k$-medoids algorithm to compute pseudo-labels ${\bm{y}}_{i}^{*}$. These are presented as colors overlayed on the eigenvectors. (d) We select the three eigenvectors whose pseudo-labels are the most “stable” with respect to several prediction models (see Section 3.2). (e) For each data feature we estimate its importance score for each of the selected eigenvectors (see Section 3.3). (f) We aggregate the feature scores across eigenvectors.
@@ 2.1 Laplacian score and representation-based feature selection
A potential drawback of the Laplacian score is its dependence on all eigenvectors, which may reduce its stability in measuring a feature’s importance to the data’s main structures. To overcome this limitation, Zhao and Liu [2007] derived an alternative score based only on a feature’s correlation to the leading Laplacian eigenvectors. A related, more sophisticated approach is Multi-Cluster Feature Selection (MCFS) [Cai et al., 2010], which computes the solutions to the generalized eigenvector problem ${\bm{L}}{\bm{v}}=\lambda{\bm{D}}{\bm{v}}$. The leading eigenvectors are then used as pseudo-labels for a regression task with $l_{1}$ regularization. Specifically, MCFS applies Least Angle Regression (LARS) [Efron et al., 2004] to obtain, for each leading eigenvector ${\bm{v}}_{i}$, a sparse vector of coefficients $\bm{\beta}^{i}\in\mathbb{R}^{p}$. The feature score is set as the maximum over the absolute values of the coefficients computed from the leading eigenvectors, $s_{j}=\max_{i}|\beta^{i}_{j}|.$ The output of MCFS is the set of features with the highest score. In the next section, we derive Spectral Self-supervised Feature Selection (SSFS), which improves upon the MCFS algorithm in several critical aspects.
@@ 3.1 Rationale
where $|A_{i}|$ denotes the size of cluster $A_{i}$. In this scenario, the zero eigenvalue of the graph Laplacian has multiplicity $k$, and the corresponding eigenvectors are equal, up to a rotation matrix, to a matrix $E\in\mathbb{R}^{n\times d}$ whose columns are equal to ${\bm{e}}^{1},\ldots,{\bm{e}}^{k}$. In this ideal setting, the $k$ leading eigenvectors are indeed suitable for use as pseudo-labels for the feature selection task. Thus, the MCFS algorithm should provide highly informative features in terms of cluster separation.
In most applications, however, cluster separation is far from perfect, and the use of leading eigenvectors may be suboptimal. Here are some common scenarios: 1) High-dimensional datasets may contain substructures in top eigenvectors, while the main structure of interest appears deeper in the spectrum. For illustration, consider the MNIST dataset visualized via t-SNE in Figure 1a. The data contains images of $3,6$ and $8$. Figure 1b shows the elements of the six leading eigenvectors of the graph Laplacian matrix, sorted by their corresponding digits. The leading eigenvector shows a clear gap between images of digit $6$ and the rest of the data. However, there is no clear separation between digits $3$ and $8$. Indeed, the next eigenvector is not associated with such a separation. Applying feature selection with this eigenvector may produce spurious features irrelevant to separating the two digits. This scenario is prevalent in the real datasets used in the experimental section. For example, Figure 2a shows four eigenvectors of a graph computed from observations containing the genetic expression data from prostate cancer patients and controls [Singh et al., 2002]. The leading two eigenvectors, however, are not associated with the patient-control separation.
@@
Figure 2: The first four Laplacian eigenvectors of two real datasets. Samples are sorted according to the class label and colored by the outcome of a one-dimensional $k$-medoids per eigenvector. The vertical bar indicates the separation between the classes. In Prostate-GE, ${\bm{v}}_{4}$ is the most informative to the class labels, and an outlier can be seen on the upper left in the third and fourth eigenvectors. In TOX-171, ${\bm{v}}_{3}$ is most informative to the class labels than ${\bm{v}}_{2}$.
@@ 3.4 Choice of Surrogate Models
8:        Compute the feature scores $s_{m}(h_{i,b})$
9:     end for
10:     for $m=1$ to $p$ do
11:        Estimate the variance of the $m$-th feature score:
@@
12:     end for
13:     $\mathcal{\hat{S}}_{i}=\sum_{m=1}^{p}\widehat{\mathrm{Var}}({s}_{m}(h_{i}))$
14:     $\mathcal{\hat{S}}\leftarrow\mathcal{\hat{S}}\cup\{\mathcal{\hat{S}}_{i}$
15:  end for
16:  Select the indices of the $k$ smallest elements in $\mathcal{\hat{S}}$ and store in $I$
17:  return  $\mathcal{Y}^{*}$, $I$
@@ 4 The importance of a proper selection of eigenvectors: analysis of the product manifold model
As described in Section 3.1, the principle of selecting the leading Laplacian eigenvectors as pseudo-labels is motivated by the case of highly separable clusters, where observations in different clusters have very low connectivity between them. In many cases, the separation between meaningful states (i.e., biological or medical conditions) may not be that clear. To illustrate this point, consider the MNIST example in Figure 1. While images with the digit $6$ constitute a cluster, the separation between digits $8$ and $3$ are not clearly separated. Figure 3a shows a scatter plot of these digits, where each image is located according to its coordinates in the third and fourth eigenvectors. Even when considering the most relevant eigenvectors, there is no clear separation between the digits. Instead, the transition between $3$ and $8$ is smooth and depends on the properties of the digits.
To provide insight into the importance of eigenvector selection, we analyze a product of manifold model. Our analysis is based on results from two research topics: (i) the convergence, under the manifold assumption, of the Laplacian eigenvectors to the eigenfunctions of the Laplace Beltrami operator associated with the manifold, and (ii) the properties of manifold products. We next provide a brief background on these two topics.
@@ 4.1 Convergence of the Laplacian eigenvectors
In many applications, high dimensional observations are assumed to reside close to some manifold $\mathcal{M}$ with low intrinsic dimensionality, which we denote by $d$. Many papers in recent decades have analyzed the relation between the Laplacian eigenvectors and the manifold structure Von Luxburg et al. [2008], Singer and Wu [2017], García Trillos et al. [2020], Wormell and Reich [2021], Dunson et al. [2021], Calder and Trillos [2022]. More formally, let ${\bm{v}}_{k}$ denote the $k$-th eigenvector of the graph Laplacian, and let $g_{k}$ denote the $k$-th eigenfunction of the Laplace-Beltrami (LB) operator. We usually assume that $g_{k}$ is normalized such that
@@
${\bm{v}}_{k}\xrightarrow[n\to\infty]{}g_{k}({\bm{X}}).$
where $g_{k}(X)$ is a vector of size $n$ containing samples of the function $g_{k}$ at the $n$ rows of the data matrix ${\bm{X}}$.
Let us consider a simple example. Suppose we have $n$ points sampled uniformly at random over the interval $[0,1]$. The LB operator over this interval is defined as the second derivative, whose eigenfunctions are the harmonic functions given by $g_{k}(x)=\cos(k\pi x)$. Figure 3 shows the three Laplacian eigenvectors computed with $n=10^{2},10^{3}$ and $3\cdot 10^{3}$ points. As $n\to\infty$, the vector $v_{k}$ converges to $g_{k}({\bm{X}})$.
Here, we use a convergence result from Cheng and Wu [2022], derived under the following assumptions: (i) The $n$ observations are generated according to a uniform distribution over the manifold, such that $\mu({\bm{x}})$ equals to a constant $\mu$. (ii) Let $\lambda_{k}$ denote the eigenvalue associated with the eigenfunction $g_{k}$. To ensure the stability of the eigenvectors, we assume a spectral gap between the smallest $K$ eigenvalues bounded away from 00 such that,
@@
(iii) The graph weights are computed by a Gaussian kernel $\exp(-\|x_{i}-x_{j}\|^{2}/\epsilon_{n})$, with a bandwidth $\epsilon_{n}\xrightarrow[n\to\infty]{}0^{+}$ that satisfies $\epsilon_{n}^{d/2+2}>C_{k}\frac{\log n}{n}$ for a constant $C_{K}$. The following result from Cheng and Wu [2022] guarantees the convergence of the Laplacian eigenvectors to samples of the LB eigenfunction.
@@
We next address the eigenvector structure in case of a product-manifold model.
@@ 4.2 The product of manifold model
For simplicity, we denote by ${\bm{v}}_{l,k}$ the $(l,k)$-th eigenvector of the Laplacian matrix, as ordered by $\lambda_{l,k}$. An example of a product of $2$ manifolds is illustrated in Figure 4b. The figure shows the leading eight eigenvectors of the graph Laplacian. The eigenvectors are indexed by the vector ${\bm{b}}=[l,k]$. The full details of this example is provided in the next section.
@@
Figure 3: Panel 3a shows a scatter plot of the noisy MNIST dataset, containing digits 3 and 8, where each image is located according to its coordinates in the third and fourth eigenvectors. Panel 3b shows the leading eigenvector of a graph computed over $n$ points on a $1D$ interval and the leading eigenfunction $\cos(\pi x)$.
4.3 Considerations for eigenvector selection in a product-manifold model